# Relaxation dynamics of integrable field theories after a global quantum quench

### v1

@@ 2.2 Local and semi-local operators
From this one can extract the semi-locality factor $l(\sigma)=-1$, as winding of the order operator around the disorder one gives a factor of $-1$. By duality, one obtains the same result in the disordered phase by exchanging the two fields. The theory possesses a global $\mathbb{Z}_{2}$ symmetry; sending $\psi\to-\psi$ is equivalent to the action of braiding the two operators. The breaking of this symmetry leads to the splitting of the Hilbert space into the Ramond (R) and the Neveu–Schwarz (NS) sectors.
@@ 4.1 Linked cluster expansion
×$\qquad\times\langle\xi_{1},-\xi_{1},\dots,\xi_{M},-\xi_{M}|-\theta_{N},\theta_{N},\dots,-\theta_{1},\theta_{1}\rangle$
@@
In the following we apply the linked cluster expansion to local and semi-local operators. The idea is to formally expand our prediction (11)  in powers of $K$, which is also contained in the rate $\Gamma$, and distinguish the terms by their time dependence [38, 42]. In fact, expanding the exponential will give terms $\propto\Gamma t$ (as well as higher powers), which originate from the annihilation poles of the form factors as shown in E. Thus in order to argue that the time evolution of local operators contains both exponential as well as power-law decay, we have to establish that (i) there exist terms $D_{M,N;\mathrm{conn}}(t)$ within $D_{M,N}(t)$ that get exponentiated, ie, the higher-order contribution $D_{M+2,N+2}(t)$ has the form $D_{M+2,N+2}(t)=-\Gamma tD_{M,N;\mathrm{conn}}(t)+\ldots$ (thus implying $f\neq 0$ [38, 42]), and (ii) that $D_{M,N}(t)$ also contains terms ${D}_{M,N;\mathrm{stat}}$ and ${D}_{M,N;\mathrm{reg}}(t)$ that do not appear with linear time dependence at higher order, ie, $D_{M+2,N+2}(t)$ does neither contain the term $-\Gamma t{D}_{M,N;\mathrm{stat}}$ nor $-\Gamma t{D}_{M,N;\mathrm{reg}}(t)$ (thus implying $g\neq 0$). Similarly, for semi-local operators we argue that for every term $D_{M,N;\mathrm{conn}}(t)$ within $D_{M,N}(t)$ that is not already growing in time, there exists a higher-order contribution $D_{M+2,N+2}(t)=-\Gamma tD_{M,N;\mathrm{conn}}(t)+\ldots$. Since in this case every term is exponentiated, it follows that $g=0$ [38, 42]. Below we focus on the diagonal contributions $D_{N,N}(t)$ to exhibit the behaviour described above.
@@ 4.1.1 Local operators
Hence the result consists of a fully connected contribution as well as a stationary one. The latter is the one-particle contribution in the LeClair–Mussardo expansion [31, 46] of the stationary value $\langle\mathcal{O}\rangle_{\rho}$. In the case of semi-local operators the latter vanishes identically.
@@ 5.1 Thermal Ising field theory
These two expectation values already contain most of the features described beforehand: the local operator $\varepsilon$ has non-vanishing overlap (and thus expectation value) with the steady state and shows power-law decay. That additional exponentially decaying terms are absent is a peculiar feature of free theories. On the contrary, the semi-local operator $\sigma$ does not relax towards a non-zero one-point expectation value and shows exclusive exponential decay, governed by the relaxation rate $\Gamma$. Both expectation values show oscillations with frequency $\sim 2m$, which is the energy required to excite a pair of excitations above the post-quench vacuum. Thus these results are consistent with (11).
@@ 5.2 Sinh-Gordon field theory
where $\mathcal{G}_{\alpha}=\left\langle 0\right|e^{\alpha g\phi}\left|0\right\rangle$ [55]. Please note that in the previous work [47] the regular piece was not considered. As compared to the energy operator in the Ising field theory, we observe exponential decay, power-law decay, and oscillations, ie, all contributions expected from (11) . The full derivation, which includes the last line, can be carried both using the linked cluster expansion or thermal form factors; it is contained in F.3.
@@ 5.3 Sine-Gordon field theory
Even thought the sine-Gordon and sinh-Gordon models are formally related by an analytic continuation in the interaction parameter, $g\to\mathrm{i}\beta$, their physical properties are rather different. The sine-Gordon model possesses stable topological excitations, named solitons and antisolitons, which classically interpolate between the minima of the cosine potential. In addition, in the attractive regime, $\beta^{2}<4\pi$, there exist stable bound states. In the quantum theory this is highlighted by the presence of poles in the scattering matrix between solitons and antisolitons. However, here we consider the repulsive regime $4\pi<\beta^{2}<8\pi$ only.
@@
$\langle e^{\mathrm{i}\beta\phi(t)}\rangle_{\psi}$ $=$ $\mathcal{G}_{\beta}+\int_{0}^{\infty}\frac{\mathrm{d}\theta}{2\pi}K^{ab}(\theta)^{*}K_{ab}(\theta)F_{1,c\mathrm{;conn}}^{\alpha}(\theta)$ (63)
@@
where the dots stand for higher order contributions in powers of $K$, and summation over the indices is implicit (and $\bar{a}=-a$). This result, as its counterpart in the sinh-Gordon theory, contains both exponential and power-law terms; both $f(t)$ and $g(t)$ scale as $t^{-3/2}$ in at late times. Thus our findings are in again line with (11).
@@
Finally, let us briefly comment on more general semi-local operators $e^{\mathrm{i}\alpha\phi}$ with $\alpha\neq\beta/2$. We first recall that a semi-classical treatment [63] yields a non-vanishing stationary expectation value $\langle e^{\mathrm{i}\alpha\phi}\rangle_{\rho}$. In contrast, the quench action approach predicts a vanishing one for the reason that $e^{\mathrm{i}\alpha\phi}$ connects different sectors in the finite volume theory; implying the vanishing of expectation values within the sectors, cf. (6). We attribute the discrepancy to the purely reflective scattering in the semi-classical approximation, which seems to prohibit relevant decay channels. We note in addition that in the attractive regime ($\beta^{2}<4\pi$) the semi-classical approach seems to miss oscillations originating from the presence of zero-momentum breathers [36, 37].
@@ Appendix B Form factors of semi-local operators in the sinh-Gordon model
In this appendix we study the twist operator $\tau$ in the sinh-Gordon model. Its defining property is that it links the different symmetry sectors associated with the $\mathbb{Z}_{2}$ symmetry $\phi\to-\phi$. Thus the twist field is semi-local and can be viewed as analog of the order and disorder operators in the Ising field theory.
@@ Appendix D Finite-volume theory
As we are not interested in the full energy spectrum at finite volume, we only need few results on the Jacobian matrix and its determinant generated by the mapping (87). We are then able to properly characterise any finite-volume state $|I_{1},\dots,I_{N}\rangle_{L}$; the initial states we discussed in C or the boundary initial state for instance (8) are then properly defined in the finite volume with the help of Bethe ansatz formalism as follows
@@ Appendix F Exponential decay and power-law behaviour from thermal form factors
the same prescription as taking the Laurent expansion for $\{\epsilon_{i}\}\to 0^{+}$ and subtracting the pole contributions. These are the only ones that have to be removed since form factors are meromorphic functions. The singular part is responsible for the exponential relaxation of the observable towards the steady state expectation values (see E), while the regular part should vanish when expectation values over the post-quench steady state are not considered. Both these results can be retrieved by means of thermal form factors as we show in this Appendix.
@@
Terms with $M_{\phi}<N$ are negligible contributions in the thermodynamics limit. We shall not consider them since they are not contributing to the expectation value.
By the means of saddle-point approximation, we can now estimate (F) in the late-time regime since the integrand in the large-volume approximation is a smooth function. We can then show that these contributions have polynomial tails at late times that dominate the exponential decay given by the annihilation poles, as we expected. Also, we can see that semi-local operators do not allow polynomial contributions, namely either the overlap $\mathcal{E}_{\phi}$ vanishes when $|\phi\rangle\in\mathrm{NS}$ or the matrix element when $|\phi\rangle\in\mathrm{R}$, as expected from (40).
@@ Appendix H Time evolution of local operators in the sine-Gordon model
$\langle\xi_{1},-\xi_{1},\dots,\xi_{N},-\xi_{N}|e^{\mathrm{i}\beta\phi}|-\rho_{N},\rho_{N},\dots,-\rho_{1},\rho_{1}\rangle$ (159)

### v2

@@ 2.2 Local and semi-local operators
From this one can extract the semi-locality factor $l(\sigma)=-1$, as winding of the order operator around the disorder one gives a factor of $-1$. By duality, one obtains the same result in the disordered phase by exchanging the two fields. The theory possesses a global $\mathbb{Z}_{2}$ symmetry; sending $\psi\to-\psi$ is equivalent to the action of braiding the two operators. The breaking of this symmetry allows to treat separatedly the Ramond (R) and the Neveu–Schwarz (NS) sectors in the evaluation of expectation values.
@@ 4.1 Linked cluster expansion
×$\qquad\times\langle\xi_{1},-\xi_{1},\dots,\xi_{N},-\xi_{N}|-\theta_{N},\theta_{N},\dots,-\theta_{1},\theta_{1}\rangle$
@@
In the following we apply the linked cluster expansion to local and semi-local operators. The idea is to formally expand our prediction (11)  in powers of $K$, which is also contained in the rate $\Gamma$, and distinguish the terms by their time dependence [38, 42]. In fact, expanding the exponential will give terms $\propto\Gamma t$ preceeded by combinatorial prefactors (as well as higher powers), which originate from the annihilation poles of the form factors as shown in E. Thus in order to argue that the time evolution of local operators contains both exponential as well as power-law decay, we have to establish that (i) there exist terms $D_{M,N;\mathrm{conn}}(t)$ within $D_{M,N}(t)$ that get exponentiated, ie, the higher-order contribution $D_{M+2,N+2}(t)$ has the form $D_{M+2,N+2}(t)\propto-\Gamma tD_{M,N;\mathrm{conn}}(t)+\ldots$ (thus implying $f\neq 0$ [38, 42]), and (ii) that $D_{M,N}(t)$ also contains terms ${D}_{M,N;\mathrm{stat}}$ and ${D}_{M,N;\mathrm{reg}}(t)$ that do not appear with linear time dependence at higher order, ie, $D_{M+2,N+2}(t)$ does neither contain the term $-\Gamma t{D}_{M,N;\mathrm{stat}}$ nor $-\Gamma t{D}_{M,N;\mathrm{reg}}(t)$ (thus implying $g\neq 0$). Similarly, for semi-local operators we argue that for every term $D_{M,N;\mathrm{conn}}(t)$ within $D_{M,N}(t)$ that is not already growing in time, there exists a higher-order contribution $D_{M+2,N+2}(t)\propto-\Gamma tD_{M,N;\mathrm{conn}}(t)+\ldots$. Since in this case every term is exponentiated, it follows that $g=0$ [38, 42]. Below we focus on the diagonal contributions $D_{N,N}(t)$ to exhibit the behaviour described above.
@@ 4.1.1 Local operators
Hence the result consists of a fully connected contribution as well as a stationary one and it agrees with the result previously found in literature for Euclidean time in a system with boundaries [47]. The latter is the one-particle contribution in the LeClair–Mussardo expansion [31, 46] of the stationary value $\langle\mathcal{O}\rangle_{\rho}$. In the case of semi-local operators the latter vanishes identically.
@@ 5.1 Thermal Ising field theory
These two expectation values already contain most of the features described beforehand: the local operator $\varepsilon$ has non-vanishing overlap (and thus expectation value) with the steady state and shows power-law decay. That additional exponentially decaying terms are absent is a peculiar feature of free theories. On the contrary, the semi-local operator $\sigma$ does not relax towards a non-zero one-point expectation value and shows exclusive exponential decay, governed by the relaxation rate $\Gamma$. Both expectation values show oscillations with frequency $\sim 2m$, which is the energy required to excite a pair of excitations above the post-quench vacuum. Thus these results are consistent with (11). Moreover, the main claim (11) agrees with expectations coming from thermal two-point functions in XY spin chain in presence of perpendicular magnetic field too [55].
@@ 5.2 Sinh-Gordon field theory
where $\mathcal{G}_{\alpha}=\left\langle 0\right|e^{\alpha g\phi}\left|0\right\rangle$ [57]. Please note that in the previous work [48] the regular piece was not considered. As compared to the energy operator in the Ising field theory, we observe exponential decay, power-law decay, and oscillations, ie, all contributions expected from (11) . The full derivation, which includes the last line, can be carried out both using the linked cluster expansion or thermal form factors; it is contained in F.3.
@@ 5.3 Sine-Gordon field theory
Even though the sine-Gordon and sinh-Gordon models are formally related by an analytic continuation in the interaction parameter, $g\to\mathrm{i}\beta$, their physical properties are rather different. The sine-Gordon model possesses stable topological excitations, named solitons and antisolitons, which classically interpolate between the minima of the cosine potential. In addition, in the attractive regime, $\beta^{2}<4\pi$, there exist stable bound states. In the quantum theory this is highlighted by the presence of poles in the scattering matrix between solitons and antisolitons. However, here we consider the repulsive regime $4\pi<\beta^{2}<8\pi$ only.
@@
$\langle e^{\mathrm{i}\beta\phi(t)}\rangle_{\psi}$ $=$ $\mathcal{G}_{\beta}+\int_{0}^{\infty}\frac{\mathrm{d}\theta}{2\pi}K^{ab}(\theta)^{*}K_{ab}(\theta)F_{1,c\mathrm{;conn}}^{\beta}(\theta)$ (63)
@@
where the dots stand for higher order contributions in powers of $K$, and summation over the indices is implicit (and $\bar{a}=-a$). This result, as its counterpart in the sinh-Gordon theory, contains both exponential and power-law terms; both $f(t)$ and $g(t)$ scale as $t^{-3/2}$ in at late times. Thus our findings are in line again with (11).
@@
Finally, let us briefly comment on more general semi-local operators $e^{\mathrm{i}\alpha\phi}$ with $\alpha\neq\beta/2$. We first recall that a semi-classical treatment [65] yields a non-vanishing stationary expectation value $\langle e^{\mathrm{i}\alpha\phi}\rangle_{\rho}$. In contrast, the quench action approach predicts a vanishing one for the reason that $e^{\mathrm{i}\alpha\phi}$ connects different sectors in the finite volume theory; implying the vanishing of expectation values within the sectors, cf. (6). We attribute the discrepancy to the purely reflective scattering in the semi-classical approximation, which seems to prohibit relevant decay channels, as explicitly shown in [66]. We note in addition that in the attractive regime ($\beta^{2}<4\pi$) the semi-classical approach seems to miss oscillations originating from the presence of zero-momentum breathers [36, 37].
@@ Appendix B Form factors of semi-local operators in the sinh-Gordon model
In this appendix we study the twist operator $\tau$ in the sinh-Gordon model. Its defining property is that it links the different symmetry sectors associated with the $\mathbb{Z}_{2}$ symmetry $\phi\to-\phi$. Thus the twist field is semi-local and can be viewed as analog of the order and disorder operators in the Ising field theory. Its form factors have been already introduced in [71] in the context of symmetry resolved entanglement of integrable quantum field theory.
@@ Appendix D Finite-volume theory
As we are not interested in the full energy spectrum at finite volume, we only need few results on the Jacobian matrix and its determinant generated by the mapping (87). We are then able to properly characterise any finite-volume state $|I_{1},\dots,I_{N}\rangle_{L}$; the initial states we discussed in C or the boundary initial state for instance (8) are then properly defined in the finite volume with the help of Bethe ansatz formalism [47] as follows
@@ Appendix F Exponential decay and power-law behaviour from thermal form factors
the same prescription as taking the Laurent expansion for $\{\epsilon_{i}\}\to 0^{+}$ and subtracting the pole contributions. These are the only ones that have to be removed since form factors are meromorphic functions. The singular part is responsible for the exponential relaxation of the observable towards the steady state expectation values (see E), while the regular part at the thermodynamic limit should tend to zero when expectation values over the post-quench steady state are not considered [77]. Both these results can be retrieved by means of thermal form factors as we show in this Appendix.
@@
Terms with $M_{\phi}<N$ are negligible contributions in the thermodynamic limit. We shall not consider them since they are not contributing to the expectation value.
By the means of saddle-point approximation, we can now estimate (F) in the late-time regime since the integrand in the large-volume approximation is a smooth function. We can then show that these contributions have polynomial tails at late times that dominate the exponential decay given by the annihilation poles, as we expected. Also, we can see that semi-local operators do not allow polynomial contributions, namely either the overlap $e^{\mathcal{E}_{\phi}}$ vanishes when $|\phi\rangle\in\mathrm{NS}$ or the matrix element when $|\phi\rangle\in\mathrm{R}$, as expected from (40).
@@ Appendix H Time evolution of local operators in the sine-Gordon model
$\langle\xi_{1},-\xi_{1},\dots,\xi_{N},-\xi_{N}|e^{\mathrm{i}\beta\phi}|-\rho_{N},\rho_{N},\dots,-\rho_{1},\rho_{1}\rangle$